# ASGO: Adaptive Structured Gradient Optimization

### v1

@@ Appendix A Details of Empirical Experiments
Table 1: Hyperparameter selection for NanoGPT experiment
Appendix B Auxiliary Lemmas
@@
These are standard trace properties. The following lemma notes the operator monotonicity of the power functions, which is a classic result [Löwner, 1934, Ando et al., 2004, Gupta et al., 2018].
@@
A proof can be found in the ICLR version of Liu et al. [2024] for Lemma G.5. ∎
Appendix C Proof of Nonsmooth Convergence of ASGO
@@
$\begin{split}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)=&{\rm tr}\left(G_{t}G_{t}^{\top}\Lambda_{t}^{-1}\right)={\rm tr}\left((\Lambda_{t}^{2}-\Lambda_{t-1}^{2})\Lambda_{t}^{-1}\right)\\ =&{\rm tr}\left(\left((\Lambda_{t}-\Lambda_{t-1})\cdot 2\Lambda_{t}-(\Lambda_{t}-\Lambda_{t-1})^{2}+\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1}\right)\Lambda_{t}^{-1}\right)\\ =&2{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right)-{\rm tr}\left((\Lambda_{t}-\Lambda_{t-1})^{2}\Lambda_{t}^{-1}\right)+{\rm tr}\left((\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1})\Lambda_{t}^{-1}\right)\\ \leq&2{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right).\end{split}$ (8)
@@
Taking $\eta=D_{\mathrm{op}}$ completes the proof.
@@
where the first equality is based on Lemma 1 and the second inequality is based on the fact that $g(x)=\sqrt{x}$ is concave for $x\geq 0$. This concludes the proof.
We also involve the following proof for the comparison between Theorem 1 and the convergence rates of Shampoo and full-matrix AdaGrad.
@@
Full-Matrix AdaGrad: $\quad\mathcal{O}\left(D_{\mathrm{F}}\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
Shampoo: $\quad\mathcal{O}\left(\sqrt{r_{G}}D_{\mathrm{F}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right)\right)$
ASGO: $\quad\mathcal{O}\left(D_{\mathrm{op}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\right)\leq\mathcal{O}\left(D_{\mathrm{op}}\cdot\sum_{j=1}^{m}\sqrt{\sum_{i=1}^{n}\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
@@
where the first inequality is based on Lemma 4 and the second inequality is simply a fact that for a vector $x$, we have $\left\|x\right\|_{2}\geq\left\|x\right\|_{1}$. Thus we prove that the proven convergence rate of ASGO is at least $D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than full-matrix AdaGrad.
@@
Therefore, the proven rate of ASGO is at least $\sqrt{r_{G}}D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than Shampoo.
Appendix D Proof of Smooth Convergence of ASGO
@@
$\leq$ $\frac{2D_{\mathrm{op}}}{T}\sqrt{\left\|L\right\|_{*}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]}+\frac{\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{T}}+\frac{\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}$
@@
Appendix E Proof of Section 6
Muon [Jordan et al., 2024] is approximately the following algorithm:
$\begin{split}B_{t}=&\mu B_{t-1}+G_{t}\\ U_{t},S_{t},V_{t}=&\text{SVD}(B_{t})\\ W_{t+1}=&W_{t}-\eta_{t}U_{t}V_{t}^{\top}\end{split}$ (10)
where $G_{t}$ is the gradient obtained at $W_{t}$ and $U_{t}\in\mathbb{R}^{m\times r_{t}},V_{t}\in\mathbb{R}^{r_{t}\times n}$ are the left and right singular vectors of $B_{t}$ with $r_{t}$ being the rank of $B_{t}$. Note that Muon implements this algorithm using the Newton-Schulz matrix iteration to approximately compute $U_{t}V_{t}^{\top}$ instead of directly employing SVD to improve computational efficiency. The relation between two original Muon forms (10) and (2) is interpreted in Bernstein and Newhouse [2024] when mentioning spectral descent. In the following discussion, we mention Muon as described in (10). We first address the equivalence of ASGO and Muon without gradient accumulation and momentum.
@@ Appendix E Proof of Section 6
If we set ASGO (Algorithm 2 version) with $\beta_{1}=\beta_{2}=0$ and $\epsilon\to 0$, $\tau=1$ and Muon (10) with $\mu=0$, ASGO is equivalent to Muon.
@@
For ASGO with $\beta_{1}=\beta_{2}=0$, the update is
$W_{t+1}=$ $W_{t}-\eta_{t}(G_{t}G_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}G_{t}$
$=$ $W_{t}-\eta_{t}(\tilde{U}_{t}\tilde{S}_{t}\tilde{S}_{t}^{\top}\tilde{U}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}\tilde{U}_{t}(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{U}_{t}^{\top}\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}\tilde{U}_{t}(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{U}_{t}^{\top}\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}\tilde{U}_{t}(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{S}_{t}\tilde{V}_{t}^{\top}$
where $\tilde{U}_{t}\in\mathbb{R}^{m\times m},\tilde{S}_{t}\in\mathbb{R}^{m\times n},\tilde{V}_{t}\in\mathbb{R}^{n\times n}$ is the full-size SVD of $G_{t}$ such that $G_{t}=\tilde{U}_{t}\tilde{S}_{t}\tilde{V}_{t}^{\top}$. We have $\tilde{S}_{t}\tilde{S}_{t}^{\top}$ is a diagonal matrix with diagonal entries being the square of singular values of $G_{t}$ and zeros. Thus we have each possibly nonzero entry of $(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{S}_{t}\in\mathbb{R}^{m\times n}$ being
$\left[(\tilde{S}_{t}\tilde{S}_{t}^{\top}+\epsilon I)^{-\frac{1}{2}}\tilde{S}_{t}\right]_{j,j}=\frac{[\tilde{S}_{t}]_{j,j}}{\sqrt{[\tilde{S}_{t}]_{j,j}^{2}+\epsilon}},\quad\text{for all $j\leq\min\left\{m,n\right\}$}$
We can see that if $[\tilde{S}_{t}]_{j,j}=0$ then the above gives 00. Thus when taking $\epsilon\to 0$, ASGO is
$W_{t+1}=W_{t}-\eta_{t}U_{t}V_{t}^{\top},$
where $U_{t}=\tilde{U}_{t}[:,:r_{t}]$ and $V_{t}=\tilde{V}_{t}[:,:r_{t}]$ with $r_{t}$ as the rank of $G_{t}$. This is the same as Muon. ∎
We then prove a nonconvex convergence result for Muon in the deterministic case with $\mu=0$.
@@
Since we assume the deterministic setting, we have $G_{t}=\nabla f(W_{t})$. Also, since we assume $\mu=0$, we have $U_{t},V_{t}$ are the left and right singular vectors of $G_{t}$ such that $G_{t}=U_{t}S_{t}V_{t}^{\top}$ Then based on the smoothness assumption, it holds that

### v2

@@ C.1 Pretraining NanoGPT
Ablation Study: Impact of Special Query/Key Design
In Appendix B, we highlighted the computational benefits of specialized processing for Query and Key matrices in Transformer architectures. To validate these theoretical advantages empirically, we conducted an ablation study comparing the performance and training stability of ASGO and Muon on pretraining NanoGPT, both with and without the specialized Query-Key processing. We maintained the optimal HP configuration established in Table 3 for all parameters except the learning rate. For each algorithm variant (with and without specialized processing), we randomly sampled learning rates from a log-uniform distribution ranging from $10^{-5}$ to $10^{-1}$. Each configuration was trained for 5 epochs on the NanoGPT model, after which we recorded the validation loss. This approach allowed us to evaluate both performance and robustness to learning rate selection. Figure 6 presents the distribution of validation losses after 5 epochs for both algorithms. The violin plots [^4] Violin plot outlines depict empirical probability density; i.e., the width of the shaded area represents the proportion of the data located there. Box plots within a violin plot display the median and inter-quartile range. reveal several key insights: (1) For ASGO (Figure 6a), the benefits of specialized Query-Key processing are substantial. The specialized implementation demonstrates a narrower, lower distribution of validation losses (centered around 2.5), indicating greater stability across different learning rates. In contrast, the vanilla implementation shows a long-tailed distribution extending beyond loss values of 6.0, with several outliers representing training instability at certain learning rates. (2) For Muon (Figure 6b), both implementations demonstrate nearly identical performance distributions, with their median validation losses and distribution shapes showing minimal differences. However, the specialized implementation achieves this comparable performance while reducing memory requirements and computational complexity. This represents a clear efficiency advantage – the specialized Query-Key processing effectively provides computational savings with no performance penalty.
(a) ASGO
(b) Muon
C.2 Pretraining GPT2
Experimental Setup:
The model configuration consists of 12 Transformer layers, 12 attention heads, and an embedding dimension of 768. The total number of training steps was 1000, with a batch size per GPU of 32. The model was trained using Distributed Data Parallel (DDP) on 4 NVIDIA V100 GPUs, employing gradient accumulation over 8 steps. To ensure a fair comparison across all optimization algorithms, a consistent learning rate schedule was utilized: a linear warmup for the first 200 steps, followed by a cosine annealing decay to a final learning rate of $1\times 10^{-5}$ for the remainder of the training process. Training was conducted on the OpenWebText dataset with a sequence length of 512 tokens.
Hyperparameter Tuning:
To ensure optimal performance for each optimizer, we conducted a hyperparameter search using grid search. The search space for learning rate (lr) and 2nd momentum $\beta_{2}$ (where applicable for the optimizer) included the following values:
  • Learning rate (lr): $[1\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},0.1,0.5]$
  • $\beta_{2}$: [0.7,0.8,0.9,0.95,0.99]
Optimal hyperparameters were selected based on the validation loss achieved after 1000 training steps. The best hyperparameter choices for pretraining GPT-2 are presented in Table 4.
Table 4: Optimal hyperparameter selection for pretraining GPT-2.
Optimizer Learning Rate $\beta_{1}$ $\beta_{2}$ Damping $\epsilon$ Update Freq. ($\tau$)
Muon 0.01 0.9 N/A N/A N/A
AdamW 0.005 0.9 0.99 $1\times 10^{-6}$ N/A
DASGO 0.01 0.9 0.99 $1\times 10^{-6}$ N/A
ASGO 0.1 0.9 0.95 $1\times 10^{-6}$ 1
C.3 Finetune GPT2-Large
Experimental Setup:
We fine-tuned the GPT-2 Large model on the WikiText-2 dataset. We utilized the GPT-2 Large model, which comprises approximately 774 million parameters. This model is a transformer-based language model pretrained on English text using a Causal Language Modeling (CLM) objective, initially developed by OpenAI and was accessed via the Hugging Face Hub. The fine-tuning was performed on the WikiText-2 dataset, a standard benchmark for evaluating language model performance. The dataset was loaded directly using standard library functions. We investigated two distinct fine-tuning objectives:
  • Causal Language Modeling (CLM): The conventional autoregressive language modeling task.
  • Fill-in-the-Middle (FIM): Adopting the methodology from Bavarian et al. (2022). This objective trains the model to infill masked text spans within a document.
All models were fine-tuned for a total of 2 epochs. For the learning rate, we employed a cosine annealing decay schedule over the course of these 2 epochs, with the learning rate decaying to 0 by the end of training. No warmup phase was used for the fine-tuning learning rate schedule. The fine-tuning process used a batch size of 16 and a sequence length of 128 tokens. For each optimizer evaluated (AdamW, Muon, ASGO, and DASGO), all hyperparameters, with the exception of the learning rate, were maintained at the same values used in our GPT-2 pretraining experiments (as detailed in Table 4).
Hyperparameter Tuning:
For the fine-tuning experiments on WikiText-2, our hyperparameter tuning focused exclusively on identifying the optimal learning rate for each optimizer under both the CLM and FIM objectives. We performed a grid search over a predefined set of learning rate candidates: $[1\times 10^{-6},5\times 10^{-5},1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3}]$. The optimal learning rate for each optimizer and fine-tuning objective combination was determined by selecting the learning rate that yielded the lowest validation perplexity on the WikiText-2 validation set after the full 2 epochs of fine-tuning. Table 5 shows the optimal learning rate.
Table 5: Optimal Learning rate from Finetuning GPT2-Large
AdamW Muon ASGO DASGO
CLM $5\times 10^{-5}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$
FIM $1\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Table 6 presents the average perplexity results obtained after fine-tuning the GPT2-Large (774M) model for 2 epochs on the WikiText-2 dataset. The $\pm$ indicates the 95% confidence intervals of the mean perplexity value over these 5 runs, starting from different random seeds.
Table 6: Finetuning GPT2-Large Perplexity
AdamW Muon ASGO DASGO
CLM 14.010 $\pm$ 0.016 13.912 $\pm$ 0.003 13.879 $\pm$ 0.010 13.841 $\pm$ 0.019
FIM 17.458 $\pm$ 2.254 15.933 $\pm$ 0.102 15.661 $\pm$ 0.097 15.451 $\pm$ 0.072
Damping parameter $\epsilon$ sensitivity study:
In this section, we discuss the practical differences between ASGO and Muon. Theoretically, as discussed in Section 6, ASGO can be viewed as an adaptive extension of Muon. Specifically, ASGO’s update rule degenerates to that of Muon in the absence of momentum and adaptive scaling (i.e., when $\beta_{1}=0$ and $\beta_{2}=0$).
Moreover, if we modify the formulas for $V_{t}$ and $W_{t+1}$ in ASGO (Algorithm 2) to:
if $\text{if }m<n\text{ then }V_{t}=\beta_{2}V_{t-1}+\left(1-\beta_{2}\right)M_{t}M_{t}^{\top}$ (4)
else $\text{else }V_{t}=\beta_{2}V_{t-1}+\left(1-\beta_{2}\right)M_{t}^{\top}M_{t}$
if $\text{if }m<n\text{ then }W_{t+1}=W_{t}-\eta_{t}(V_{t}+\epsilon I)^{-\frac{1}{2}}M_{t}$
else $\text{else }W_{t+1}=W_{t}-\eta_{t}M_{t}(V_{t}+\epsilon I)^{-\frac{1}{2}}$
and set $\beta_{2}=0$ ASGO becomes identical to Muon, even with momentum. However, in practical training implementations, a subtle yet crucial difference exists between ASGO and Muon. This distinction arises because ASGO applies an additional preconditioner with $\epsilon$ damping to the momentum term, a step not present in the standard Muon update.
To precisely study these differences, we set up an experiment using the Muon optimizer as the background optimizer to train a GPT2-model on the NanoGPT dataset for 1000 steps. At each step, we computed and compared the Muon and ASGO update directions $\Delta W$ in the embedding, query, key, value, and MLP layers. Specifically, we computed:
  • Standard Muon Update Direction (SVD-based): $\Delta W^{\text{Muon}}_{\text{SVD}}=U_{M}V_{M}^{T}$, where $U_{M}$ and $V_{M}$ are the left and right singular vectors obtained from the first-order momentum $M$.
  • Modified ASGO Update Direction (SVD-based, $\epsilon=0$): $\Delta W^{\text{ASGO}}_{\text{SVD},\epsilon=0}=V_{t}^{-\frac{1}{2}}M_{t}$, where $V_{t}^{-\frac{1}{2}}$ is computed via standard Singular Value Decomposition (SVD).
  • Modified ASGO Update Direction (SVD-based, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{SVD},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed via standard SVD.
  • Modified ASGO Update Direction (Coupled Newton, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{CN},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed using the Coupled Newton algorithm for 50 iteration steps. Higham [2008], Shi et al. [2023]
  • Modified ASGO Update Direction (Newton-Schulz, $\epsilon=10^{-8}$): $\Delta W^{\text{ASGO}}_{\text{NS},\epsilon=10^{-8}}=(V_{t}+10^{-8}I)^{-\frac{1}{2}}M_{t}$, where $(V_{t}+10^{-8}I)^{-\frac{1}{2}}$ is computed using the Newton-Schulz algorithm for 50 iteration steps. [^5] The Newton-Schulz algorithm is derived from Algorithm 6.35 on Page 153 in Higham [2008]. We utilized the quintic version with parameters $(a,b,c)=(2,-1.5,0.5)$ as described in Jordan et al. [2024]. Higham [2008], Bernstein and Newhouse [2024b], Jordan et al. [2024]
[…]
(a) $\operatorname{CosineSimilarity}(\Delta W^{\text{Muon}}_{\text{SVD}},\Delta W^{\text{ASGO}}_{\text{SVD},\epsilon=0})$
(b) $\operatorname{CosineSimilarity}(\Delta W^{\text{Muon}}_{\text{SVD}},\Delta W^{\text{ASGO}}_{\text{SVD},\epsilon=10^{-8}})$
(c) $\operatorname{CosineSimilarity}(\Delta W^{\text{Muon}}_{\text{SVD}},\Delta W^{\text{ASGO}}_{\text{CN},\epsilon=10^{-8}})$
(d) $\operatorname{CosineSimilarity}(\Delta W^{\text{Muon}}_{\text{SVD}},\Delta W^{\text{ASGO}}_{\text{NS},\epsilon=10^{-8}})$
Figure 7: Comparison of Cosine Similarities (CSs) between the Muon update direction ($\Delta W^{\text{Muon}}_{\text{SVD}}$) and various modified ASGO update directions in different layers (MLP, Value, Key, Query, Embedding) during 1000 steps of GPT2 pretraining. Subfigures (a) and (b) plot CSs for SVD-based ASGO update directions with $\epsilon=0$ and $\epsilon=10^{-8}$, respectively. Subfigure (c) and (d) plot CSs for Coupled Newton-based and Newton-Schulz-based, respectively, ASGO update directions with $\epsilon=10^{-8}$.
Unlike AdamW, ASGO exhibits sensitivity to the choice of the $\epsilon$ damping parameter. As depicted in Figure 7(a), when no damping term is used ($\epsilon=0$), ASGO’s update direction, though exhibiting some instability, largely reconstructs Muon’s update direction, with cosine similarities generally exceeding 0.8, particularly for the MLP and Embedding layers, as training progresses. However, the introduction of a small damping hyperparameter ($\epsilon=10^{-8}$) significantly alters these similarities. Figure 7(b) shows that even with precise SVD computation, the similarities for layers like Value and Key only reach approximately 0.7 gradually. This suggests that the small $\epsilon$ value can introduce and accumulate errors during training, potentially leading to a degradation in optimizer performance, a behavior not typically observed in other common adaptive methods like AdamW. This phenomenon highlights the critical importance of small singular values in the preconditioning process, as applying a damping term can disproportionately affect them, thereby altering the matrix multiplication and amplifying errors. To circumvent this sensitivity, we opted to use an $\epsilon$ value of 0 in our large-scale GPT2 pretraining experiments described in Section 7.2. In this setting, we directly compute the SVD of the preconditioning matrix and its inverse square root. The results presented in Figure 2 indeed demonstrate that ASGO achieves performance highly similar to Muon.
In Figure 7(c) and (d), we explore the use of more computationally efficient methods, Coupled Newton (CN) and Newton-Schulz (NS), respectively, to approximately compute $V_{t}^{-\frac{1}{2}}$ as alternatives to SVD. The results show that, aside from some numerical stability issues observed in the Embedding layer for both CN and NS at certain steps—which is reasonable given that the sparse nature of embedding layers often requiring larger damping parameters—the cosine similarities for the remaining parameters are highly comparable to those obtained with SVD. This demonstrates the feasibility of employing efficient matrix algorithms to replace SVD, thereby enhancing ASGO’s computational efficiency. This direction represents one of our promising avenues for future research.
Appendix D Auxiliary Lemmas for the Proof
@@
The following lemma notes the operator monotonicity of the power functions, which is a classic result [Löwner, 1934, Ando et al., 2004, Gupta et al., 2018].
@@
A proof can be found in Lemma G.5 of the ICLR version of Liu et al. [2024]. ∎
Appendix E Proof of Nonsmooth Convergence of ASGO
@@
$\begin{split}{\rm tr}\left(G_{t}^{\top}\Lambda_{t}^{-1}G_{t}\right)=&{\rm tr}\left(G_{t}G_{t}^{\top}\Lambda_{t}^{-1}\right)={\rm tr}\left((\Lambda_{t}^{2}-\Lambda_{t-1}^{2})\Lambda_{t}^{-1}\right)\\ =&{\rm tr}\left(\left((\Lambda_{t}-\Lambda_{t-1})\cdot 2\Lambda_{t}-(\Lambda_{t}-\Lambda_{t-1})^{2}+\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1}\right)\Lambda_{t}^{-1}\right)\\ =&2{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right)-{\rm tr}\left((\Lambda_{t}-\Lambda_{t-1})^{2}\Lambda_{t}^{-1}\right)+{\rm tr}\left((\Lambda_{t-1}\Lambda_{t}-\Lambda_{t}\Lambda_{t-1})\Lambda_{t}^{-1}\right)\\ \leq&2{\rm tr}\left(\Lambda_{t}-\Lambda_{t-1}\right),\end{split}$ (9)
@@
Taking $\eta=D_{\mathrm{op}}$ completes the proof. ∎
@@
where the first equality is based on Lemma 1 and the second inequality is based on the fact that $g(x)=\sqrt{x}$ is concave for $x\geq 0$. This concludes the proof. ∎
We also provide the following proof for the comparison between Theorem 1 and the convergence rates of Shampoo and full-matrix AdaGrad.
@@
[…]
ASGO: $\quad\mathcal{O}\left(D_{\mathrm{op}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{2}}\right)\right)\leq\mathcal{O}\left(D_{\mathrm{op}}\cdot\sum_{j=1}^{m}\sqrt{\sum_{i=1}^{n}\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right).$
@@
where the first inequality is based on Lemma 4 and the second inequality is simply a fact that for any vector $x$, we have $\left\|x\right\|_{2}\geq\left\|x\right\|_{1}$. Thus we prove that the proven convergence rate of ASGO is at least $D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than full-matrix AdaGrad.
@@
Therefore, the proven rate of ASGO is at least $\sqrt{r_{G}}D_{\mathrm{F}}/D_{\mathrm{op}}$ times faster than Shampoo. ∎
Appendix F Proof of Smooth Convergence of ASGO
@@
$\leq$ $\frac{2D_{\mathrm{op}}}{T}\sqrt{\left\|L\right\|_{*}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]}+\frac{\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{T}}+\frac{2\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T}.$
@@
Appendix G Proof of Section 6
Muon [Jordan et al., 2024] is essentially the following algorithm:
$\begin{split}B_{t}=&\ \mu B_{t-1}+G_{t}\\ U_{t},S_{t},V_{t}=&\ \text{Compact SVD}(B_{t})\\ W_{t+1}=&\ W_{t}-\eta_{t}U_{t}V_{t}^{\top},\end{split}$ (11)
where $G_{t}$ is the gradient obtained at $W_{t}$ and $U_{t}\in\mathbb{R}^{m\times r_{t}},V_{t}\in\mathbb{R}^{r_{t}\times n}$ are the matrices of the $r_{t}$ left and right singular vectors corresponding to the non-zero singular values of $B_{t}$ that are the diagonal components of the diagonal matrix $S_{t}\in\mathbb{R}^{r_{t}\times r_{t}}$ with $r_{t}$ being the rank of $B_{t}$. Note that $U_{t}S_{t}V_{t}^{\top}$ is referred to as the "compact" or "Reduced" SVD of $B_{t}$. Muon implements this algorithm using Newton-Schulz matrix iterations to approximately compute $U_{t}V_{t}^{\top}$ instead of directly employing SVD to improve computational efficiency. The relationship between (11), with $B_{t}=G_{t}$, and (2) is interpreted in Bernstein and Newhouse [2024b] (see Story II). In the following discussion, we refer Muon as the one described in (11). We first show that ASGO and Muon, without gradient accumulation and momentum, are equivalent.
@@ Appendix G Proof of Section 6
If in ASGO (Algorithm 2 version) we set $\beta_{1}=\beta_{2}=0,\epsilon=0$, and $\tau=1$, and in Muon (11) we set $\mu=0$, ASGO is equivalent to Muon.
@@
Let $U_{t}S_{t}V_{t}^{\top}$ be the compact SVD of $G_{t}$. For ASGO, the update is
$W_{t+1}=$ $W_{t}-\eta_{t}(G_{t}G_{t}^{\top})^{-\frac{1}{2}}G_{t}$
$=$ $W_{t}-\eta_{t}({U}_{t}{S}_{t}{S}_{t}^{\top}{U}_{t}^{\top})^{-\frac{1}{2}}{U}_{t}{S}_{t}{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}{U}_{t}({S}_{t}^{2})^{-\frac{1}{2}}{U}_{t}^{\top}{U}_{t}{S}_{t}{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}{U}_{t}{S}_{t}^{-1}{S}_{t}{V}_{t}^{\top}$
$=$ $W_{t}-\eta_{t}{U}_{t}{V}_{t}^{\top}.$
Clearly, this is the same as Muon. ∎
We now prove a nonconvex convergence result for Muon in the deterministic case with $\mu=0$.
@@
Since we assume the deterministic setting, we have $G_{t}=\nabla f(W_{t})$. Also, since we assume $\mu=0$, we have $U_{t},V_{t}$ are the left and right singular vectors of $G_{t}$ such that $G_{t}=U_{t}S_{t}V_{t}^{\top}$. Then based on the smoothness assumption, it holds that